# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v3

@@ 5.1 Knowledge and Semantic Information
Consider mapping $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
$\mathcal{E}_{C}(A)=A\cup\{k\in\mathcal{K}\mid\exists f\in\mathcal{F}_{C}:k=f(A)\},$ (7)
where $\mathcal{F}_{C}$ is the set of all computational transformations with cost (complexity) lower than or equal to $C$ and $k=f(A)$ denotes knowledge $k$ derivable from $A$ using transformation $f\in\mathcal{F}_{C}$. We call mapping $\mathcal{E}_{C}$ the emergence operator if it satisfies the following expansion rules:
$\mathcal{E}_{C}(\mathcal{E}_{C}(A))=\mathcal{E}^{(2)}_{C}(A)\supseteq\mathcal{E}_{C}(A)\quad\text{and}\quad\lim_{n\to\infty}\mathcal{E}_{C}^{(n)}(A)=A^{*}.$ (8)
For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process with emergence operator:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{{satisfies}}\quad\mu_{C}(\theta(t+1))\geq\mu_{C}(\theta(t)).$ (9)
Furthermore, with unlimited computation, reasoning creates no new information:
$\mu_{\infty}(\theta(t+1))=\mu_{\infty}(\theta(t)).$ (10)
The definitions above clearly require rigorous refinements, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination. In particular, they are meant to lay ground for the principle of semantic information conservation and its potential violations. As we will see, that idea is closely related to the following property of response.
Definition 5 (Bounded Response Creativity).
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Response $r$ is creatively bounded if and only if:
$\mu_{C}(K(r))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))).$ (11)
We address the settings in which a large subset of knowledge encoded in LLM can be redundant, so it doesn’t necessarily contain more information than its components. Clearly, that assumption is violated when LLM combines knowledge elements to fabricate facts or produce emergent insights. Then,
@@
which means novel insights come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?
We view reasoning as a transformation that reveals information already present in the knowledge representation. After the reasoning, certain patterns may become evident that were computationally intractable to observe in the original representation. However, we stand at the position that reasoning does not create new information.
To see that, consider a similarity relation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues. Matrices $D$ and $A$ contain the same information, but it is much easier to understand the fundamental properties of $A$ when we represent it as $D$ in the basis of eigenvectors in matrix $V$. The properties of $A$ has always been encoded in $A$, but they become accessible when we transform $A$ into $D$.
Resolution of the Uncertainty Resolution Paradox
If reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable. But that means the reasoning must have created new information, contrary to what we have assumed. It seems we have a paradox.
@@
Notice that for all practical purposes the information measurement $\mu_{C}(X,q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time). Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X,q)\leq\mu_{C}(\mathcal{E}_{C}(X),q)$. Notice, that is possible only with organized information providing knowledge for reasoning.
That idea above is thus described by the reasoning dynamics:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t),q).$ (12)
Bounded reasoning makes hidden knowledge accessible within a computational budget. So, when that budget is infinite, reasoning creates no new information, i.e.,
@@
To illustrate that argument, consider the following example:
[…]
@@
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0)),$
@@ 5.2 Statements, Queries, and Ground Truth
$T(q)=\{k\in\mathcal{K}\mid k\text{ is {relevant} to }K(q)\text{ and {labeled} as true}\}.$ (14)
@@ 5.3 Auction-Theoretic Model of LLM Inference
The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $-\nabla_{\theta_{i}}L_{W}$).
@@ 5.5 Properties of Hallucination-Free Mechanisms
There are four fundamental properties that a hallucination-free response mechanism should ideally satisfy. We can interpret them as constraints that should be incorporated in model training. We characterize them formally referring to the notation of fundamental results of mechanism design and auction theory.
@@
Theorem 1 (Information–Contribution and Bounded Creativity).
Let $K_{0}=K(g(\mathbf{0}))$ be the knowledge in the query-only response and let
@@
be the additional knowledge agent $i=1,\dots,n$ can reveal to support full response $g(s)$. Let us also define the information contribution of agent $i$ as:
@@
If $K_{i}\cap K_{j}=\varnothing$ for $i\neq j$, then
$\sum_{i=1}^{n}p_{i}(s,\theta)=\mu_{C}\bigl{(}K(g(s))\bigr{)}-\mu_{C}\bigl{(}K_{0}\bigr{)}$ (22)
and $\sum_{i}p_{i}=0$ if and only if $\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$.
@@
Since knowledge increments $\{K_{i}\}_{i=1}^{n}$ are pairwise disjoint, so that $K_{i}\perp_{C}K_{j}$, it follows from the semantic measure properties that:
$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})$ (23)
and similarly,
$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j\neq i}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j}).$ (24)
Therefore, the marginal contribution of agent $i$ simplifies to:
$p_{i}=\bigl{[}\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})\bigr{]}-\bigl{[}\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j})\bigr{]}=\mu_{C}(K_{i}).$ (25)
@@
$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$ (26)
Then, by assumption, the full response satisfies $K(g(s))=K_{0}\cup\bigcup_{j=1}^{n}K_{j}$. We conclude that
$\sum_{i}p_{i}=0\quad\text{if and only if}\quad\mu_{C}(K(g(s)))=\mu_{C}(K_{0}),$ (27)
which completes the proof. ∎
The result above reveals an intriguing consequence of the semantic information conservation principle. Namely, if the semantic information is indeed conserved, then
$\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$ (28)
means the response contains no more semantic information than the baseline query-only response. Therefore, the response is not introducing new information, which means it is of bounded creativity from Definition 5. So, in such a case, LLM cannot genuinely create new knowledge, only reorganize existing information. Apparent creativity in LLMs may actually be nothing more than sophisticated recombination of existing patterns. Theorem 1 quantifies this limitation precisely.
@@
The information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge. In our dynamical systems framework, this corresponds to the constraint that the measured output cannot exceed what can be produced from the initial state plus what becomes accessible through state transitions within computational budget $C$.
Notice that the principle does not prevent redistribution of information. It is permitted as long as the information transfers perfectly balance. Indeed, since $p_{i}=p_{i}(s,\theta)$ incorporates the strategic profile, it admits strategies withholding information. Within the information conservation regime we do not require that strategies utilize all available knowledge states. Inference process can withhold information, as it often does in practice.
@@
Then the information conservation rule is satisfied when $p$ counts exchanged bits of knowledge. Namely, $p_{1}(s,\theta)=1$ for providing $A$, $p_{2}(s,\theta)=2$ for providing $\{D,E\}$, and $p_{3}(s,\theta)=-3$ for absorbing information into $\{A,D,E\}$. We have

### v4

@@ 5.1.3 Limits of Knowledge and Reasoning
Let $\mathcal{F}_{C}$ be a complete set of all computational transformations with cost (complexity) lower than or equal to $C$, such that $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ if $C<D$ and such that for any $g,h\in\mathcal{F}_{\infty}=\bigcup_{C}\mathcal{F}_{C}$ there exists $h\in\mathcal{F}_{\infty}$ for which $g(A\cup f(A))=h(A)$. Emergence operator is a mapping $\mathcal{E}_{C}\colon\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
$\mathcal{E}_{C}(A)=A\cup\{B\in\mathcal{K}\mid B=f(A)\ \text{for some monotone}\ f\in\mathcal{F}_{C}\}.$ (3)
Furthermore, to represent reasoning with unlimited computational budget, we define:
$\mathcal{E}_{\infty}(A)=A\cup\{B\in\mathcal{K}\mid B=f(A)\ \text{for some monotone}\ f\in\mathcal{F}_{\infty}\}.$ (4)
Finally, for $n\geq 0$ we have:
$\mathcal{E}_{C}^{(0)}=\texttt{id}\quad\text{and}\quad\mathcal{E}_{C}^{(n+1)}=\mathcal{E}_{C}\circ\mathcal{E}_{C}^{(n)}.$ (5)
Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest. When $A\subseteq\mathcal{E}_{C}(A)$, the discovered additional elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible. That aligns with the pragmatist or constructivist view that understanding involves active construction rather than passive reception.
When $\mathcal{E}_{C}(A)=A\cup B$, we say that $B$ emerges from reasoning based on $A$. Then $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$. In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information as expected. Whenever $A\subseteq A\cup B$, we have
$\mu_{C}(A)\leq\mu_{C}(A\cup B)=\mu_{C}(\mathcal{E}_{C}(A))\leq\mu_{C}(A)+\mu_{C}(B).$ (6)
The emergence operator is monotonic, which means the reasoning process it describes is also monotonic. Whenever $A\subseteq B$, we must have $\mathcal{E}_{C}(A)\subseteq\mathcal{E}_{C}(B)$. Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A)\subseteq\mathcal{E}_{D}(A)$. By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
Theorem 2 (Fixed-point of Reasoning).
For any $A\subseteq\mathcal{P}(\mathcal{K})$ and budget $C>0$, there exists a fixed point of the emergence operator $\mathcal{E}_{C}$:
$A^{*}_{C}=\bigcup_{n\geq 0}^{|\mathcal{K}|}\mathcal{E}_{C}^{(n)}(A)=\mathcal{E}_{C}(A^{*}_{C}).$ (7)
Proof.
Since $(\mathcal{P}(\mathcal{K}),\subseteq)$ is a complete lattice and $\mathcal{E}_{C}$ is monotone, by the Knaster-Tarski fixed point theorem Tarski1955ALF there exists a fixed point $A^{*}_{C}$ of reasoning with $\mathcal{E}_{C}$. Finiteness of $\mathcal{K}$ ensures convergence within $|\mathcal{K}|$ steps. ∎
Let us apply the results above to show how $\mu_{\infty}$ can be constructed with $\mathcal{E}_{\infty}$. One example exploits the reasoning operations as follows:
$\mu_{\infty}(A)=\log|\mathcal{K}|\cdot|\bigcup_{n}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A)|.$ (8)
That $\mu_{\infty}$ satisfies axioms demanded in Definition 3. Clearly, $\mu_{\infty}$ is bounded and
$\mu_{\infty}(\emptyset)=\log|\mathcal{K}|\cdot|\bigcup_{n}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(\emptyset)|=\log|\mathcal{K}|\cdot|\emptyset|=0.$
If $A\subseteq B$, then $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(B)$ implies $\mu_{\infty}(A)\leq\mu_{\infty}(B)$. When $A\perp_{\infty}B$, no element of $A^{*}_{\infty}$ can be derived from $B$ (and vice versa). Since $A\cap B=\emptyset$, we conclude $A^{*}_{\infty}\cap B^{*}_{\infty}=\emptyset$ and $|(A\cup B)^{*}_{\infty}|=|A^{*}_{\infty}|+|B^{*}_{\infty}|$. That means $\mu_{\infty}(A\cup B)=\mu_{\infty}(A)+\mu_{\infty}(B)$ for computationally independent sets. That shows the proposed formula satisfies all required properties of semantic information measure.
All that brings us to the following conclusion. With infinite computing and reasoning power, total information encoded in $A$ must be preserved. Indeed, that bold (philosophical) claim is at the center of our investigations. Let us then justify it rigorously.
Definition 6 (Equivalence of Knowledge).
We say that $X\subseteq\mathcal{K}$ and $Y\subseteq\mathcal{K}$ are equivalent, $X\equiv Y$, if and only if $X\subseteq\mathcal{E}_{\infty}(Y)$ and $Y\subseteq\mathcal{E}_{\infty}(X)$.
Notice that equivalence of knowledge does not mean $A$ is a fixed point if emergence, i.e., $A=\mathcal{E}_{\infty}(A)$ does not mean $A\equiv\mathcal{E}_{\infty}(A)$. If we have $A=\mathcal{E}_{\infty}(A)$, then nothing more can be derived from $A$ by any reasoning rule from $A$. That statement is very strong and is not true when $A\equiv\mathcal{E}_{\infty}(A)$. Suppose $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ allows for getting $y=f(x)$. Then $\mathcal{E}_{\infty}(A)=\{x,y\}\supset\{x\}=A$, but $A\equiv\mathcal{E}_{\infty}(A)$.
To proceed we will also need the idempotence of reasoning with unlimited computational budget.
Theorem 3 (Idempotence of unlimited reasoning).
For any $A\subseteq\mathcal{K}$ we have:
$\mathcal{E}_{\infty}(A)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A)).$ (9)
Proof.
By definition $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$. To show $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))\subseteq\mathcal{E}_{\infty}(A)$, we must find $h\in\mathcal{F}_{\infty}$ such that $h(A)=g(A\cup\bigcup f_{i}(A))$ for $g,f_{i}\in\mathcal{F}_{\infty}$. But since we can always find such $h\in\mathcal{F}_{\infty}$, we conclude that any conclusion $g(\mathcal{E}_{\infty}(A))$ derived from $\mathcal{E}_{\infty}(A)$ by a reasoning rule $g\in\mathcal{F}_{\infty}$ can be also derived directly from $A$ as $h(A)=g(A\cup\bigcup f_{i}(A))$. Therefore, $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))\subseteq\mathcal{E}_{\infty}(A)$ as required. ∎
With all the building blocks available, we can now to prove that unlimited reasoning ($C=\infty$) can reveal but never create information. The semantic content of a knowledge base is invariant under unlimited reasoning of emergence operator. Apparent information gain at finite computational budget emerges as an effect of limited computational access.
Theorem 4 (Information preservation).
For any $A\subseteq\mathcal{K}$, if $\mu_{\infty}$ is constant on equivalent sets of knowledge, i.e., $\mu_{\infty}(X)=\mu_{\infty}(Y)$ whenever $X\equiv Y$, then:
$\mu_{\infty}(\mathcal{E}_{\infty}(A))=\mu_{\infty}(A).$ (10)
Proof.
We must show equivalence of knowledge encoded in $A$ and $\mathcal{E}_{\infty}(A)$, i.e., we demand $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$. But, by Theorem 3, we see that $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))=\mathcal{E}_{\infty}(A)$ $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))=\mathcal{E}_{\infty}(A)$ as requested. Since $A\equiv\mathcal{E}_{\infty}(A)$ and $\mu_{\infty}$ is constant on equivalent sets of knowledge, we conclude that $\mathcal{E}_{\infty}(A)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$. ∎
The Uncertainty Resolution Paradox and its Resolution
We address the general settings in which any subset of knowledge encoded in LLM can be redundant, so
$\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B).$
But when LLM combines knowledge elements to fabricate facts or produce emergent insights, it seems that
@@
so novel insights may come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?
First, subadditivity of semantic information measure establishes a baseline where uncontrolled synergy of knowledge cannot appear. We remove from our analytical toolkit the potential measurement bias of information surplus that any information aggregation mechanism might generate.
Second, we prove that hallucinations are inevitable even when knowledge is subadditive. In that setting, by Theorem 4, reasoning does not create new information. It is a transformation that reveals information already present in any knowledge representation. Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination. However, we should take care of a paradox those claims apparently bring into existence.
Consider a similarity transformation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues. That transformation is an example of reasoning function. Matrices $D$ and $A$ contain the same information, but it is much easier to understand the fundamental properties of $A$ when we represent it as $D$ in the basis of eigenvectors in matrix $V$. The properties of $A$ has always been encoded in $A$, but they become accessible when we transform $A$ into $D$. So, after the reasoning, certain patterns become evident that were impossible to observe in the original representation.
But, if reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable. Doesn’t it mean the reasoning must have created new information, contrary to what we have just claimed? It seems we have a paradox.
@@
Notice that for all practical purposes the information measurement $\mu_{C}(X)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time). Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X)\leq\mu_{C}(\mathcal{E}_{C}(X))$. That is possible only with organized information providing knowledge for reasoning. These are precisely the statements of Theorem 1 and Theorem 4.
That idea describes reasoning as a dynamical system:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t))$ (11)
with evolving knowledge state. Bounded reasoning makes hidden knowledge accessible within a computational budget. When that budget is infinite, reasoning creates no new information, i.e.,
@@
To illustrate that argument, consider the following learning trajectory:
$\theta(0)$ $=[1,1,1,0,0]\text{ (initial knowledge state)},$
$\theta(1)$ $=\mathcal{E}_{C}(\theta(0))=[1,1,1,1,1]\text{ (state after reasoning)},$
@@
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0))$
@@
5.1.4 Knowledge Representation in LLMs
The above Knowledge Representation finds its direct counterpart in the structure and state of a transformer architecture of LLMs.
The abstract knowledge space $\mathcal{K}$ is the high-dimensional embedding space of the model. Every fact, concept, and linguistic pattern the model has learned is represented by a specific element of that space. The metric $d_{\mathcal{K}}$ is typically implemented as cosine similarity or Euclidean distance, where proximity between vectors signifies a closer semantic relationship. Model parameters $\mathcal{W}$ correspond to the full set of trained weights in the transformer, including the attention matrices ($W_{Q},W_{K},W_{V},W_{O}$) and the matrices of the MLP layers. The mapping $\Phi(W)$ represents the totality of information, patterns, and factual associations stored within these trained weights. Activation patterns $\mathcal{A}$ correspond to the state of the residual stream during a single forward pass for a given input. An activation pattern is therefore a specific trajectory through the network’s computational graph. Finally, the mapping $\Psi(a_{i})$ represents the information made computationally accessible during a specific inference step and encoded in the residual stream.
The process of information processing under computational constraints maps to the sequential forward pass transformer inference.
The idea of computational budget $C$ can be represented by the depth or size of the network, e.g.,the number of transformer blocks. A deeper network has a larger budget $C$, allowing for more reasoning steps. In the context of multi-step reasoning, $C$ can also represent the number of iterations allowed in a Chain-of-Thought (CoT) sequence.
The semantic information measure $\mu_{C}$ can be linked to the model’s objective function. For a given input $A$, $\mu_{C}(A)$ can be understood as the measurable reduction in cross-entropy loss that the model achieves by processing $A$ with a budget $C$.
Finally, the emergence operator $\mathcal{E}_{C}$ can be identified with a transformer block (one multi-head attention layer followed by one MLP layer). This block applies a computational transformation $\mathcal{E}_{C}$ to the residual stream at layer $t$, producing the updated state at layer $t+1$, making information that was latent in the input explicitly accessible.
@@ 5.2 Statements, Queries, and Ground Truth
[…]
@@ 5.3 Auction-Theoretic Model of LLM Inference
The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $\nabla_{\theta_{i}}L_{W}$).
@@ 5.5 Properties of Hallucination-Free Mechanisms
There are four fundamental properties that a hallucination-free response mechanism should ideally satisfy. We can interpret them as constraints that should be incorporated in model training. We characterize them formally referring to the fundamental results we have established earlier.
5.5.1 Truthfulness
The fundamental requirement for a reliable inference system is that its components do not misrepresent their knowledge. The principle of truthfulness, also known as incentive compatibility in mechanism design, ensures that each agent’s optimal strategy is to accurately represent its privately held informations.
@@ 5.5.1 Truthfulness
5.5.2 Semantic Information Conservation
Second, we want to consider mechanisms that cannot create knowledge ex nihilo. The Semantic Information Conservation principle demands that the net informational contribution across all agents sums to zero. This property is the formal guardrail against ungrounded fabrication and ensures that any generated response, no matter how creative, is ultimately derived from the system’s existing knowledge
@@ 5.5.2 Semantic Information Conservation
Definition 13 (Bounded Response Creativity).
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q))$. Response $r$ is creatively bounded if and only if:
$\mu_{C}(K(r))\leq\mu_{C}(B^{*}_{C}).$ (19)
Theorem 5 (Bounded Creativity of Chain of Thoughts).
Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q))$. Then, let us assume that the knowledge in the query-only response meets the following condition:
$K_{0}=K(g(\mathbf{0}))\subseteq B^{*}_{C}.$ (20)
Next, let
@@
be independent and privately held knowledge that each agent $i=1,\dots,n$ can reveal, such that:
$K_{i}\perp_{C}K_{j}\text{ for all }i\neq j\quad\text{and}\quad K_{i}\perp_{C}K_{0}\text{ for all }i.$ (22)
Let us also define the information contribution of agent $i$ as:
@@
Assume the response mechanism implements the following chain-of-thought processes:
$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\right),$ (24)
such that for null-strategy responses we have:
$K_{0}\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)).$ (25)
If the semantic information is conserved, $\sum_{i=1}^{n}p_{i}=0$, then the response is creatively bounded, i.e.,
$\mu_{C}(K(r))\leq\mu_{C}(B^{*}_{C}).$ (26)
@@
Since $K_{0},K_{1},\ldots,K_{n}$ are pairwise computationally independent, by semantic measure satisfies additivity
$p_{i}(s,\theta)=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})-\mu_{C}(K_{0})-\sum_{j\neq i}\mu_{C}(K_{j})=\mu_{C}(K_{i}).$
@@
$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$
Therefore, when $\sum_{i=1}^{n}p_{i}(s,\theta)=0$, we have $\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})$. By the independence assumption, $K_{i}\perp_{C}K_{j}\text{ for all }i\neq j$, and semantic information measure non-negativity, with $p_{i}=\mu_{C}(K_{i})\geq 0$,
$\mu_{C}(K_{i})=0\text{ for all }i=1,\ldots,n.$
Then, the nullset and positivity property of $\mu_{C}$ together with independence condition $K_{i}\perp_{C}K_{0}$ imply $K_{i}=\varnothing$ for all $i=1,\ldots,n$.
By the response generation principle:
$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\right)=\mathcal{E}_{C}(K_{0}\cup\varnothing)=\mathcal{E}_{C}(K_{0}).$
By assumption, $K_{0}\subseteq B^{*}_{C}$. Applying the emergence operator to both sides, we preserve the inclusion due to monotonicyty of $\mathcal{E}_{C}$. Therefore, by the definition of a fixed point (Theorem 2), we have:
$\mathcal{E}_{C}(K_{0})\subseteq\mathcal{E}_{C}(B^{*}_{C})=B^{*}_{C}.$
Combining the inclusions yields:
$K(g(s))\subseteq\mathcal{E}_{C}(K_{0})\subseteq\mathcal{E}_{C}(B^{*}_{C})=B^{*}_{C}.$
Finally, applying the monotonicity of the semantic information measure $\mu_{C}$:
$\mu_{C}(K(g(s)))\leq\mu_{C}(B^{*}_{C}).$
Theorem 5 captures direct correspondence between the conservation principle and bounded creativity under the idealization of disjoint information contributions. Namely, if the semantic information is indeed conserved, the response contains no more semantic information than the baseline query-only response. That is true for both standard inference and chain-of-thought (CoT) reasoning.
In standard reasoning, the response mechanism $g$ applies $\mathcal{E}_{C}$ through forward passes in the neural network. With CoT reasoning, multi-step explicit reasoning is iterative applications of reasoning rules within the computational budget $C$. In both cases, the emergence operator $\mathcal{E}_{C}$ bounds what can be derived. When information is conserved, even sophisticated CoT reasoning cannot create genuinely new information. It can only reorganize and make explicit what was already latent in the baseline knowledge $K_{0}$. This explains why CoT improves performance on complex reasoning tasks (by better utilizing $\mathcal{E}_{C}$) but cannot overcome the fundamental impossibility of simultaneous truthfulness, conservation, revelation, and optimality.
@@
Furthermore, the information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge. In our dynamical systems framework, this corresponds to the constraint that the measured output cannot exceed what can be produced from the initial state plus what becomes accessible through state transitions within computational budget $C$.
However, it is important to notice that the principle does not prevent redistribution of information. It is permitted as long as the information transfers perfectly balance. Indeed, since $p_{i}=p_{i}(s,\theta)$ incorporates the strategic profile, it admits strategies withholding information. Within the information conservation regime we do not require that strategies utilize all available knowledge states. Inference process can withhold information, as it often does in practice.
@@
Suppose elements $\{B,C\}$ represent contextual information about unrelated topics. When $s_{1}(\theta_{1})=A$, the agent optimally reveals only relevant knowledge, satisfying both truthfulness and conservation. Had $\{B,C\}$ been relevant but withheld, the mechanism would violate the property of Relevant Knowledge Revelation. Then the information conservation rule is satisfied when $p$ counts exchanged bits of knowledge. Namely, $p_{1}(s,\theta)=1$ for providing $A$, $p_{2}(s,\theta)=2$ for providing $\{D,E\}$, and $p_{3}(s,\theta)=-3$ for absorbing information into $\{A,D,E\}$. We have
@@
5.5.3 Relevant Knowledge Revelation
Third, to be useful, a mechanism must not only prevent the fabrication of knowledge but also prevent the omission of relevant truth. The principle of Relevant Knowledge Revelation serves as a participation constraint, guaranteeing that any agent possessing knowledge valuable for the query contributes. This ensures the model fully utilizes its available information rather than abstaining or being overly cautious.